OpenStudy (anonymous):

Evaluate g(3) if g(x) = 2x + 1.

OpenStudy (a_clan):

2(3) +1 =7

OpenStudy (anonymous):

\[g(3) = 7\]

OpenStudy (anonymous):

\[g(x) = 2x + 1\] \[g(3) = 2(3) + 1\] \[g(3) = 7\]

OpenStudy (llort):

if x=3, you replace the x with a 3... g(3)= 2*3+1 that will give you...

OpenStudy (anonymous):

2(3)+1=7